# Demo: Soccer Information Retrieval via Natural Queries using SoccerRAG

### v1

@@ I Introduction
The burgeoning interest in Retrieval Augmented Generation (RAG), fueled by the rapid advancements in Large Language Models (LLMs), has paved new avenues for exploring innovative use cases across various domains, notably within multimodal information retrieval [1]. Despite the wide applicability of RAG frameworks in enhancing the capabilities of generative AI for open-domain question answering and beyond, its potential in the sports domain, particularly within soccer analytics, remains largely untapped. Our recent research has endeavored to bridge this gap by harnessing multimodal datasets inherent to the soccer broadcast pipeline and integrating RAG with different data modalities. We have presented SoccerRAG, a framework for retrieving multimodal soccer information using natural language queries, from an augmented soccer dataset based on SoccerNet [2, 3], which includes game videos with image frames and audio, timestamped captions (transcribed audio), annotations for game events, and player information.
In this paper, we demonstrate the use of the SoccerRAG framework through the command-line interface (CLI) and a novel user interface (UI). Figure 1 presents an overview of the SoccerRAG framework, with the core components (database, feature extractor, feature validator, and SQL agent) as presented earlier, wrapped by a novel interactive UI which serves as a user-friendly tool for visual interactions. The UI was created using the Chainlit framework. Chainlit is an open-source Python package that enables developers to build production-ready conversational AI applications, providing features for quick integration. It offers integrations with popular libraries and frameworks, such as OpenAI and LangChain, and allows for custom front-ends with React-based user interfaces [4]. The SoccerRAG integrates functionality for data representation, feature extraction and validation, database querying (based on our proposed database schema for the augmented version of the SoccerNet dataset), and a novel extractor-validator chain. The open source implementation for SoccerRAG is accessible under [5] and fully reproducible through the instructions provided therein, which are detailed below.
@@ II-B Artifacts
Figure 2 presents the structure of the SoccerRAG codebase, publicly accessible under [5]. The data folder is used by the core pipeline to read in the source dataset (including the original SoccerNet data, as well as our augmented league and team tables), and storing the database files. The src folder contains the configuration and executable files. The root directory includes the main executable (main.py), environment variables (.env), and a list of dependencies (requirements.txt). A proof-of-concept deployment of the SoccerRAG UI can be found under: https://simulamet-host-soccerrag.hf.space/
G SoccerRAG
? data/
{ dataset/
0 augmented_leagues.csv
0 augmented_teams.csv
<downloaded assets>
/ src/
/ database.py
/ extractor.py
/ sql_chain.py
2 conf/
schema.json
sqls.json
.env
/ app.py
/ main.py
/ main_cli.py
requirements.txt
/ setup.py
Figure 2: Structure of the SoccerRAG codebase.
@@ II-C Requirements
The SoccerRAG codebase requires Python 3.12 or above to be run. The requirements.txt file in the codebase specifies the required packages, which can be installed through pip. SoccerRAG does not require a GPU, and can be run on any machine with a CPU. For the experiments presented in the research paper, a machine with a Windows 11 operating system, Intel(R) Core(TM) i5-9300H CPU @2.40GHz, and 16GB memory has been used. SoccerRAG requires an OpenAI API key, which should be specified in the .env file. In order to log all queries to the OpenAI endpoints, it is possible to set a LangSmith API key [6]. This makes it possible to monitor the cost of each call, and keep history.
@@ II-D1 Configuration
TABLE I: SoccerRAG configuration parameters (M: mandatory).
@@ II-D2 Running via UI
Figure 3 presents an example screenshot, depicting alternative background colors [^2] SoccerRAG UI background color is configurable: click the user icon in the top right corner, select settings, then toggle the dark mode switch. for the UI, as well as the intermediate feedback to users that the SoccerRAG provides for each operational step, which are retrieved from the feature extractor and feature validator components in the core pipeline. The meaningful response returned to the user prompt "Can you calculate Real Madrids [sic] home advantage for the 2015/16 season?" demonstrates that the framework is able to grasp soccer-related concepts such as a home advantage (which refers to the difference in scores achieved by a team in home vs. away games).
@@
Figures 4 and 5 present further example screenshots, demonstrating the capabilities of SoccerRAG which range from database-wide statistics (example 2: "How many yellow cards did messi [sic] get in the 2015-16 season on home turf?" is returned with the correct aggregation across all games of the desired season), to supporting multi-format information presentation (example 3: "List all games played by ManU [sic] in the 16-17 season. / Give the result as a markdown table with following format / HomeTeam AwayTeam Score Venue Attendance Date").
@@ III Conclusion
We demonstrate SoccerRAG, a framework that leverages RAG and LLMs to efficiently retrieve multimodal soccer information via natural language queries, enhancing the accessibility of sports datasets by allowing for intuitive user interactions with complex data archives. In addition to the core pipeline presented in our research paper, we have introduced a novel UI to facilitate visual user interactions with the framework in a chatbot-like manner. The UI serves as a wrapper around the core pipeline, whereby it passes the user query to the feature extractor component, after which the query is validated by the feature validator against the database, the query with validated properties are passed to SQL agent, which retrieves data from the database and returns a response, and the UI presents the response to the user.
We present the entire codebase for the SoccerRAG as open source software, along with instructions to run the framework either via CLI or via UI. Our experiments up to now demonstrate SoccerRAG’s capability to accurately interpret complex queries and facilitate dynamic user engagements with soccer content. We believe that SoccerRAG exemplifies the transformative potential of AI in sports analytics, promising enriched global fan experiences and advances in the accessibility of sports information.

### v2

@@ Abstract
Index Terms:
association football, information retrieval, large language models, natural language processing, sports, UI
@@ I Introduction
The burgeoning interest in Retrieval Augmented Generation (RAG), fueled by the rapid advancements in Large Language Models(LLMs), has paved new avenues for exploring innovative use cases across various domains, notably within multimodal information retrieval [1]. Despite the wide applicability of RAG frameworks in enhancing the capabilities of generative AI for open-domain question answering and beyond, its potential in the sports domain, particularly within soccer analytics, remains largely untapped. Our recent research has endeavored to bridge this gap by harnessing multimodal datasets inherent to the soccer broadcast pipeline and integrating RAG with different data modalities. We have presented SoccerRAG [2], a framework for retrieving multimodal soccer information using natural language queries, from an augmented soccer dataset based on SoccerNet [3, 4], which includes game videos with image frames and audio, timestamped captions (transcribed audio), annotations for game events, and player information.
In this paper, we demonstrate the use of the SoccerRAG framework through a user interface (UI) based on ChainLit [5], as well as the command-line interface (CLI). Figure 1 presents an overview of the SoccerRAG framework, with the core components (database, feature extractor, feature validator, and SQL agent) as presented earlier, wrapped by a novel interactive UI which serves as a user-friendly tool for visual interactions. The UI was created using the Chainlit framework. Chainlit is an open-source Python package that enables developers to build production-ready conversational AI applications, providing features for quick integration. It offers integrations with popular libraries and frameworks, such as OpenAI and LangChain, and allows for custom front-ends with React-based user interfaces [5]. The SoccerRAG integrates functionality for data representation, feature extraction and validation, database querying (based on our proposed database schema for the augmented version of the SoccerNet dataset), and a novel extractor-validator chain. The open source implementation for SoccerRAG is accessible under [6] and fully reproducible through the instructions provided therein, which are detailed below.
@@ II-B Artifacts
The SoccerRAG codebase is publicly accessible under [6]. The data folder is used by the core pipeline to read in the source dataset (including the original SoccerNet data, as well as our augmented league and team tables), and storing the database files. The src folder contains the configuration and executable files. The root directory includes the main executable (main.py), environment variables (.env), and a list of dependencies (requirements.txt). A proof-of-concept deployment of the SoccerRAG UI can be found under: https://simulamet-host-soccerrag.hf.space/
@@ II-C Requirements
The SoccerRAG codebase requires Python 3.12 or above to be run. The requirements.txt file in the codebase specifies the required packages, which can be installed through pip. SoccerRAG does not require a GPU, and can be run on any machine with a CPU. For the experiments presented in [2], a machine with a Windows 11 operating system, Intel(R) Core(TM) i5-9300H CPU @2.40GHz, and 16GB memory has been used. SoccerRAG requires an OpenAI API key, which should be specified in the .env file. In order to log all queries to the OpenAI endpoints, it is possible to set a LangSmith API key [7]. This makes it possible to monitor the cost of each call, and keep history.
@@ II-D1 Configuration
TABLE I: SoccerRAG parameters (M: mandatory).
@@ II-D2 Running via UI
Figure 2 presents an example screenshot [^2] SoccerRAG UI background color is configurable: click the user icon in the top right corner, select settings, then toggle the dark mode switch. depicting the intermediate feedback to users that the SoccerRAG provides for each operational step, which are retrieved from the feature extractor and feature validator components in the core pipeline. The meaningful response returned to the user prompt "Can you calculate Real Madrids [sic]333sic (adverb) indicates that the quoted matter is exactly as found in the original source, including erroneous spelling, punctuation, or grammar. home advantage for the 2015/16 season?" demonstrates that the framework is able to grasp soccer-related concepts such as a home advantage (which refers to the difference in scores achieved by a team in home vs. away games).
@@
Figures 3 and 4 present further example screenshots, demonstrating the capabilities of SoccerRAG which range from database-wide statistics (example 2: "How many yellow cards did messi [sic] get in the 2015-16 season on home turf?" is returned with the correct aggregation across all games of the desired season), to supporting multi-format information presentation (example 3: "List all games played by ManU [sic] in the 16-17 season. / Give the result as a markdown table with following format / HomeTeam AwayTeam Score Venue Attendance Date").
@@ II-D3 Running via CLI
II-D4 Results
In [2], we demonstrate that this framework can improve the hit-rate of the answers from 20% to 80%.
@@ III Conclusion
We demonstrate SoccerRAG, a framework that leverages RAG and LLMs to efficiently retrieve multimodal soccer information via natural language queries, enhancing the accessibility of sports datasets by allowing for intuitive user interactions with complex data archives. In addition to the core pipeline presented in [2], we introduce a novel UI to facilitate visual user interactions with the framework in a chatbot-like manner. The UI serves as a wrapper around the core pipeline, whereby it passes the user query to the feature extractor component, after which the query is validated by the feature validator against the database, the query with validated properties are passed to SQL agent, which retrieves data from the database and returns a response, and the UI presents the response to the user. Our experiments up to now demonstrate SoccerRAG’s capability to accurately interpret complex queries and facilitate dynamic user engagements with soccer content.
We present the entire codebase for the SoccerRAG as open source software, along with instructions to run the framework either via CLI or via UI. The underlying framework concept can be used for other applications as well, provided that a source dataset and SQL schema are available. Looking forward, SoccerRAG is poised for enhancements such as real-time processing and integration with emerging AI technologies, which will broaden its application beyond soccer, or even sports, and improve user experience.